OpenStudy (anonymous):

answer correctly and you get a cookie and a medal and best friend forever who is going to fail if he doesn't get these right

OpenStudy (anonymous):

Jess plans to increase the amount of money she saves each month. She can increase her savings in the following ways: Month 1 2 3 4 5 Option 1 20 22 24.2 26.62 29.28 Option 2 20 22.2 24.4 26.6 28.8 Option 3 20 24 28.8 34.56 41.47 If Jess wants to increase her savings exponentially, which option(s) should she choose? Only option 1, because it shows equal increases in equal intervals of time Only option 2, because it shows equal increases in equal intervals of time Either option 1 or option 3, because they show increases in time by the same percentages Either option 2 or option 3, because they show increases in time by the same percentages Question 9 (Multiple Choice Worth 4 points) (04.05) The functions f(x) and g(x) are described using the following equation and table: f(x) = 3(1.02)x x g(x) -1 -4 0 6 1 8 2 10 Which statement best compares the y-intercepts of f(x) and g(x)? The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus y-intercept of f(x). Question 10 (Multiple Choice Worth 4 points) (05.05) Determine the type of boundary line and shading for the graph of the inequality 3x + y greater than or equal to 10. Dashed line with shading on the side that includes the origin Solid line with shading on the side that does not include the origin Dashed line with shading on the side that does not include the origin Solid line with shading on the side that includes the origin Question 11 (Multiple Choice Worth 4 points) (04.01) What is 3 to the power of 3 over 2 equal to? cube root of 9 square root of 9 cube root of 27 square root of 27 Question 12 (Multiple Choice Worth 4 points) (04.03) The expression 600(1.004274)4x shows the quarterly interest rate Markey earns after x years. Which expression shows the x isolated in the exponent in an equivalent form? 600(4.017096)x 600(0.251069)x 600(1.001067)x 600(1.017206)x Question 13 (Multiple Choice Worth 4 points) (05.02) The table below shows four systems of equations: System 1 System 2 System 3 System 4 4x − 5y = 2 3x − y = 8 4x − 5y = 2 10x − 7y = 18 4x − 5y = 2 3x − 8y = 4 4x − 5y = 2 10x + 3y = 15 Which pair of systems will have the same solution? System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to two times the second equation in system 1 System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to two times the second equation in system 2 System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to three times the second equation in system 1 System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to three times the second equation in system 2 Question 14 (Multiple Choice Worth 4 points) (04.05) A function is shown below: f(x) = (1.93)x What does the function represent? Exponential growth of 7% Exponential decay of 7% Exponential growth of 93% Exponential decay of 93% Question 15 (Multiple Choice Worth 4 points) (04.05) The tenth, eleventh, and twelfth terms of a sequence are shown in the table below: Term number 10 11 12 Term 42 47 52 Which of the following shows the first five terms of the sequence? 3, -2, -7, -12, -17 2, 7, 12, 17, 22 -3, 2, 7, 12, 17 -2, -7, -12, -17, -22 Question 16 (Multiple Choice Worth 4 points) (04.05) Heidi is preparing for the national gymnastics competition. The table below shows the number of hours Heidi spent preparing for a gymnatics competition over a period of five months: Month 1 2 3 4 5 Hours 2 3.5 5 6.5 8 Did Heidi increase the number of hours of practice linearly or exponentially? Linearly, because the table shows an equal increase in hours for an equal increase in months Exponentially, because the table shows an equal increase in hours for an equal increase in months Linearly, because the table shows that hours increase by an equal factor for an equal increase in months Exponentially, because the table shows that hours increase by an equal factor for an equal increase in months